# denotes equal contribution

### v1

@@ Abstract
Recent advancements in Vision-Language (VL) models have sparked interest in their deployment on edge devices, yet challenges in handling diverse visual modalities, manual annotation, and computational constraints remain. We introduce EdgeVL, a novel framework that bridges this gap by seamlessly integrating dual-modality knowledge distillation and quantization-aware contrastive learning. This approach enables the adaptation of large VL models, like CLIP, for efficient use with both RGB and non-RGB images on resource-limited devices without the need for manual annotations. EdgeVL not only transfers visual language alignment capabilities to compact models but also maintains feature quality post-quantization, significantly enhancing open-vocabulary classification performance across various visual modalities. Our work represents the first systematic effort to adapt large VL models for edge deployment, showcasing up to $15.4\%$ accuracy improvements on multiple datasets and up to 93-fold reduction in model size.
@@ 1 Introduction
  • We highlight EdgeVL’s gain in accuracy across multiple datasets and detail its efficiency improvements on diverse GPU tiers. Our code will be made publicly available.
@@ 2.2 Cross-modality Knowledge Distillation
Knowledge distillation facilitates the transfer of expertise between models through various strategies such as response distillation [21], feature distillation [40], and relation distillation [6]. In the realm of cross-modality knowledge distillation, methods like those presented by [22] and [16] transfer knowledge by sharing weights between models of different modalities. [45] employs a strategy where multiple student models are encouraged to align their predictions with a teacher model, enhancing cross-modal action recognition. CMKD [23] introduces a combination of feature and response distillation to propagate knowledge from the LiDAR modality to the RGB modality. Unlike these methods, which are tailored for labeled datasets, our work uniquely addresses the challenge of knowledge distillation using unlabeled pairs of RGB and non-RGB images.
@@ 3.4.1 QAT Meets Contrastive Learning
To explore the impact of model quantization on the expressiveness of features, we initially apply PTQ [26] on $\varPhi_{{img}}^{{stu}}$ and examine the discriminative nature of features post-quantization. As demonstrated in Fig. 3, the discriminability of features notably declines after quantization, leading to a misalignment between text and image representations. This reduction in feature clarity compared to the full-precision model underscores the necessity of employing QAT for the final image encoder optimization. QAT enhances the model by incorporating fake quantization during training, which simulates quantization effects through quantization-aware matrix multiplication, followed by finetuning the pre-trained model to adapt to these effects (see supplementary for details of QAT). A pivotal aspect of implementing QAT is selecting an appropriate loss function that maintains or even improves the discriminative capability of features within the quantization framework. Traditional knowledge distillation loss, as formulated in Eq. 4, aims to align the student model’s features with those of a pre-trained teacher model. However, it might not exploit the full potential of quantized models in achieving robust and discriminative features. In this context, we propose the integration of contrastive learning loss, which is designed to cultivate representations resilient to minor perturbations and enhances the separation in the feature space between similar and dissimilar instances. This approach is predicated on the robustness of contrastive learning, which should support the acquisition of invariant representations less affected by the distortions due to quantization. Fig. 3 illustrates how employing contrastive learning loss in conjunction with QAT not only mitigates the discriminative power reduction but potentially increases the discriminability of the student encoder after quantization.
@@ 4.2.4 Efficiency
Tab. 2 presents the efficiency profile of EdgeVL, incorporating two key metrics: (i) the size of EdgeVL’s final image encoder, and (ii) the encoder’s latency/throughput across various GPU tiers. Notably, EdgeVL significantly reduces the model size compared to the original image encoder in CLIP-G and CLIP-B. To CLIP-G, the reduction is up to $\sim 93$-fold for Swin-T. To CLIP-B whose accuracy is much worse, the size reduction ratio is of $\sim 6$-fold and $\sim 4$-fold with the two student backbones. Furthermore, inference latency is halved on two embedded GPUs (AGX and Nano), while throughput on high-end GPUs like RTX4090, sees increases of approximately 42% and 93% for Swin-T and ViT-S backbones, respectively. These improvements make EdgeVL suitable for edge devices under resource constraints.
@@ 4.3.1 Quantization Aware Contrastive Learning
Tab. 3 demonstrates the open-vocabulary classification accuracy of int8 quantized models using different methods for edge device deployment, focusing on static quantization results, though dynamic quantization aligns similarly and is detailed in the supplementary. PTQ significantly lowers accuracy, whereas QAT with stage-1 loss Eq. 4 improves it but doesn’t reach the EdgeVL benchmark, sometimes lagging by over $9.7\%$ (40.3% vs. 50.0%). Unlike conventional methods, EdgeVL consistently maintains high accuracy across various student models post-quantization, highlighting its effectiveness and robustness in preserving model performance after quantization, outperforming other strategies including QViT and PTQ by significant margins.
@@ 4.3.2 Dual-modality Knowledge Distillation
To assess the effect of dual-modality knowledge distillation, we evaluate the outcomes of our stage-1 training against CMKD(non-RGB)[23] and CMKD(RGB) [23], which resemble our stage-1 but rely on single-modality training. Tab. 4 illustrates the accuracy of these methods using the DAT-T model on two datasets. Our findings show that dual-modality training draws on the strengths of RGB-modality and non-RGB-modality training, yielding high accuracy for both RGB and non-RGB modalities, whether non-RGB are depth images from ScanNet or infrared images from EuroSAT. On average, our dual-modality training method achieves an accuracy increase of $15.0\%$ on ScanNet and $13.1\%$ on EuroSAT. Notably, the accuracy improvements on non-RGB modality (e.g., 38.6% vs. 37.8% and 61.5% vs. 61.2%) lead us to hypothesize that RGB images function as a type of data augmentation for non-RGB images in dual-modality training.
@@ 4.3.3 Cutoff Ratio of $\tau_{c}$
Tab. 7 presents the accuracy of EdgeVL (Swin-T) on EuroSAT. It indicates that $\tau_{c}=0.25$ achieves the highest accuracy compared to the other cutoff ratios. A very small $\tau_{c}$ fails to fully train the student model, whereas a very large $\tau_{c}$ introduces more noisy samples, thus diminishing the effectiveness of knowledge transfer.
@@
Table 6: Effect of different sampling strategies
Table 7: Effect of different training strategies
$\tau_{c}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Sampling [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Training [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 6: Effect of different sampling strategies
Table 7: Effect of different training strategies
@@ 4.4 Cross-Dataset Performance
We evaluated the generalization performance of models adapted by EdgeVL on ScanNet and further tested their open-vocabulary classification on the unseen NYU2 and SUNRGBD datasets, which have distinct label sets: i. SUNRGBD dataset [43] contains 5,285 training and 5,050 test RGBD images with 19 scene categories. The RGBD images are captured by multiple RGBD sensors, such as Kinect, Asus Xtion, and Intel RealSense; and ii. NYU2 dataset [42] contains 795 training and 654 test RGBD images with 10 scene categories, which are captured by a Microsoft Kinect sensor.
Results As shown in Tab. 8, a trade-off is revealed: while EdgeVL-enhanced encoders significantly improved depth image accuracy (up to 25.4% with DAT-T), there was a slight decrease in RGB image performance compared to the pre-trained CLIP model’s ViT-G encoder. This was expected due to the substantial reduction in model size post-adaptation/quantization (e.g., Swin-T’s 56MB vs. CLIP-G’s 5,213MB. Considering CLIP’s training involved 400 million image-text pairs, and EdgeVL adapted using only 4,725 RGB and depth image pairs from ScanNet, it suggests that the limited adaptation dataset size may also affect the model’s generalization ability.

### v2

# Self-Adapting Large Visual-Language Models to Edge Devices across Visual Modalities

@@ Abstract
Recent advancements in Vision-Language (VL) models have sparked interest in their deployment on edge devices, yet challenges in handling diverse visual modalities, manual annotation, and computational constraints remain. We introduce EdgeVL, a novel framework that bridges this gap by seamlessly integrating dual-modality knowledge distillation and quantization-aware contrastive learning. This approach enables the adaptation of large VL models, like CLIP, for efficient use with both RGB and non-RGB images on resource-limited devices without the need for manual annotations. EdgeVL not only transfers visual language alignment capabilities to compact models but also maintains feature quality post-quantization, significantly enhancing open-vocabulary classification performance across various visual modalities. Our work represents the first systematic effort to adapt large VL models for edge deployment, showcasing up to $15.4\%$ accuracy improvements on multiple datasets and up to 93-fold reduction in model size. Code available at https://github.com/ramdrop/edgevl.
@@ 1 Introduction
  • We highlight EdgeVL’s gain in accuracy across multiple datasets and detail its efficiency improvements on diverse GPU tiers.
@@ 2.2 Cross-modality Knowledge Distillation
Knowledge distillation facilitates the transfer of expertise between models through various strategies such as response distillation [22], feature distillation [41], and relation distillation [7]. In the realm of cross-modality knowledge distillation, methods like those presented by [23] and [17] transfer knowledge by sharing weights between models of different modalities. [46] employs a strategy where multiple student models are encouraged to align their predictions with a teacher model, enhancing cross-modal action recognition. CMKD [24] introduces a combination of feature and response distillation to propagate knowledge from the LiDAR modality to the RGB modality. Recently, CLIP has also been explored for transferring knowledge from 2D images to 3D scenes [54, 3, 51]. Unlike these methods tailored for labeled datasets, our work uniquely addresses the challenge of knowledge distillation using unlabeled pairs of RGB and non-RGB images.
@@ 3.4.1 QAT Meets Contrastive Learning
To explore the impact of model quantization on the expressiveness of features, we initially apply PTQ [27] on $\varPhi_{{img}}^{{stu}}$ and examine the discriminative nature of features post-quantization. As demonstrated in Fig. 3, the discriminability of features notably declines after quantization, leading to a misalignment between text and image representations. This reduction in feature clarity compared to the full-precision model underscores the necessity of employing QAT for the final image encoder optimization. QAT enhances the model by incorporating fake quantization during training, which simulates quantization effects through quantization-aware matrix multiplication, followed by finetuning the pre-trained model to adapt to these effects (see supplementary for details of QAT). A pivotal aspect of implementing QAT is selecting an appropriate loss function that maintains or even improves the discriminative capability of features within the quantization framework. Traditional knowledge distillation loss, as formulated in Eq. 4, aims to align the student model’s features with those of a pre-trained teacher model. However, it might not exploit the full potential of quantized models in achieving robust and discriminative features. In this context, we propose the integration of contrastive learning loss, which is designed to cultivate representations resilient to non-discriminative features and enhances the separation in the feature space between similar and dissimilar instances. This approach is predicated on the robustness of contrastive learning, which should support the acquisition of invariant representations less affected by the distortions due to quantization. Fig. 3 illustrates how employing contrastive learning loss in conjunction with QAT not only mitigates the discriminative power reduction but potentially increases the discriminability of the student encoder after quantization.
@@ 4.3.1 Quantization Aware Contrastive Learning
Tab. 3 demonstrates the open-vocabulary classification accuracy of int8 quantized models using different methods for edge device deployment, focusing on static quantization results, though dynamic quantization aligns similarly and is detailed in the supplementary. PTQ significantly lowers accuracy, whereas QAT with stage-1 loss Eq. 4 improves it but doesn’t reach the EdgeVL benchmark, sometimes lagging by over $9.7\%$ (40.3% vs. 50.0%). In comparison, stage-2 in EdgeVL introduces contrastive learning in the quantization-aware training process, which improves the student model’s discriminability and accuracy despite quantization. The result highlights EdgeVL’s effectiveness and robustness in preserving model performance after quantization, outperforming other strategies including QViT and PTQ by significant margins.
@@ 4.3.2 Dual-modality Knowledge Distillation
We evaluate the outcomes of our stage-1 training against CMKD(non-RGB)[24] and CMKD(RGB) [24], which resemble our stage-1 but rely on single-modality training. Tab. 4 illustrates the accuracy of these methods using the DAT-T model. Our findings show that dual-modality training draws on the strengths of RGB-modality and non-RGB-modality training, yielding high accuracy for both modalities, whether non-RGB are depth images from ScanNet or infrared images from EuroSAT. Our dual-modality training method achieves an increase of $15.0\%$ on ScanNet and $13.1\%$ on EuroSAT in average accuracy. Notably, the accuracy boosts on non-RGB modality (e.g., 38.6% vs. 37.8% and 61.5% vs. 61.2%) lead us to hypothesize that RGB images function as a type of data augmentation for non-RGB images in dual-modality training.
@@ 4.3.3 Cutoff Ratio of $\tau_{c}$
Tab. 7 presents the accuracy of EdgeVL (Swin-T) on EuroSAT, where $\tau_{c}=0.25$ brings the highest accuracy compared to the other cutoff ratios. A very small $\tau_{c}$ fails to fully train the model, whereas a very large $\tau_{c}$ introduces noisy samples, thus diminishing the effectiveness of knowledge transfer.
@@
$\tau_{c}$ Refer to caption Refer to caption Refer to caption
@@
Table 6: Effect of different sampling strategies
Sampling Refer to caption Refer to caption Refer to caption
@@
Table 7: Effect of different training strategies
Training Refer to caption Refer to caption Refer to caption
@@ 4.4 Cross-Dataset Performance
We assessed the generalization performance of EdgeVL by training the models on ScanNet and evaluating their open-vocabulary classification accuracy on the unseen SUNRGBD and NYU2 datasets: i. SUNRGBD dataset [44] contains 5,285 training and 5,050 test RGBD images with 19 scene categories. The images are captured by multiple RGBD sensors, including Kinect, Asus Xtion, and Intel RealSense; and ii. NYU2 dataset [43] contains 795 training and 654 test RGBD images with 10 scene categories, which are captured by a Microsoft Kinect sensor.
As shown in Tab. 8, a trade-off is revealed: while EdgeVL-enhanced encoders significantly improved depth image accuracy (up to 25.4% with DAT-T), there was a slight decrease in RGB image accuracy compared to the pre-trained CLIP model’s ViT-G encoder. This was expected due to the substantial reduction in model size post-quantization (e.g., Swin-T’s 56MB vs. CLIP-G’s 5,213MB). Besides, CLIP’s training involved 400 million image-text pairs, while EdgeVL used only 4,725 RGB and depth image pairs from ScanNet, suggesting that the limited adaptation dataset size may affect the model’s generalization ability.